OpenStudy (anonymous):

Factor completely: 4x2 – 121 (2x – 11)(2x + 11) (2x – 11)(2x – 11) (4x – 121)(x + 1 ) (4x – 121)(x – 1)

Parth (parthkohli):

DIFFERENCE OF SQUARES :P

Parth (parthkohli):

\( \color{Black}{\Rightarrow a^2 - b^2 = ( a + b)(a - b) }\)

OpenStudy (anonymous):

SO I USE THAT TO SLOVE IT

OpenStudy (anonymous):

SOLVE

Parth (parthkohli):

Exactly.

OpenStudy (anonymous):

CAN YOU SHOW ME HOW TO DO IT

OpenStudy (studyhard):

I can help :)

OpenStudy (anonymous):

THANK YOU

OpenStudy (studyhard):

a^2-b^2=(a-b)(a+b). Therefore, 4x^2-121=(2x-11)(2x+11). You just need to square root the a^2 and b^2, then fit it into the (a-b)(a+b). Hope you get what I mean! :) No prob!

OpenStudy (anonymous):

Factor completely: 16x2 – 49 (16x – 49)(x + 1) (16x – 49)(x – 1) (4x – 7)(4x + 7) (4x – 7)(4x – 7)

OpenStudy (studyhard):

Same concept. Try this one yourself?

OpenStudy (anonymous):

SO ALL I HAVE TO DO IS 16X^2-49

OpenStudy (studyhard):

16x^2 is a^2. So a will be 4x. Then 49 is b^2. So b is 7. Therefore, (a-b)(a+b) will be (4x-7)(4x+7)

OpenStudy (anonymous):

IM GOING TO PUT THE NEXT ONE BUT I WANT TO SEE IF I CAN GET THE ANSWER BY MYSELF

OpenStudy (anonymous):

Factor completely: 18x2 – 98 2(3x – 7)(3x –7) 2(3x + 7)(3x – 7) (6x – 14)(3x – 7) (6x + 14)(3x + 7)

OpenStudy (anonymous):

I DONT HAVE A CALCULATOR BUT IS IT THE SECOND ONE

OpenStudy (studyhard):

YES!!! YOU GOT IT! Awesome! Continue practicin these questions and soon you'll be more familiar! :)